# Invisible Ears at Your Fingertips: Acoustic Eavesdropping via Mouse Sensors

### v1

@@ V-A Targeted Mouse Sensors
We focus on two state-of-the-art mice sensors currently available in consumer-grade mice – Paw3395 and Paw3399 – both produced by the manufacturer PixArt Imaging Inc. [49, 50]. These sensors are in the Razer Viper 8KHz [8] and Darmoshark M3-4KHz [7], respectively. An overview of the mice sensors technology used in our testing methodology is provided in Table I. We also list similar mice that are vulnerable to the same class of exploit as the Razer and Darmoshark mice in Appendix -B. These additional mice share at least one of the following attributes: (i) a Paw3395 or Paw3399 sensor or (ii) a poll rate at or above 4KHz.
@@ V-C Mouse Data Preprocessing
To address the nonuniformity in the collected mouse signals, we apply a signal processing pipeline that includes sequential resampling and filtering stages. We use a sinc-based resampling technique that allows efficient and accurate interpolation and reconstruction of fragmented, nonuniform, and noisy signals [51]. This resampling stage is equivalent to applying a modified Whittaker–Shannon interpolation formula:
@@ V-D Waveform Signal Filtering
It is important to note that environmental and sensor noise are a point of concern in our data collection scheme. Therefore, we implement a custom-tuned Wiener Filter adapted to the sensor noise introduced by having a high DPI setting. We collect 10-minute segments of the mouse jitter which are used to estimate the noise profile, and we couple the noise spectrum with the true spectrum of speech which we acquire from our selected datasets (see Sec V-E3). By applying a Wiener filter, the attacker can efficiently increase the Signal-to-Noise Ratio (SNR) of the resulting waveform [52, 53].
@@ V-E2 Speech Keyword Classification Models
This task involves the automated extraction of significant keywords from reconstructed audio waveforms. Formally, it is defined as a mapping function that translates the predicted audio waveform signal, denoted as $\hat{Y}$, into a set of key terms $K_{1},K_{2},\ldots,K_{n}$} that encapsulate critical information within the speech. We use wav2vec2[58], designed to extract and classify keywords. Wav2vec2 is based on a transformer architecture, enabling a large temporal receptive field that spans 25 ms of audio. The Wav2vec2 model is trained on a supervised dataset, for 20 epochs each, with a batch size of 64 and Adam optimizer.
@@ V-E3 Training Datasets
The ML models incorporated in the Mic-E-Mouse pipeline are trained using English voice data from existing speech recognition datasets, including (i) the CSTR VCTK Corpus (version 0.92) dataset, which contains approximately 54 hours of human speech recorded from 110 English speakers [59], and (ii) the AudioMNIST dataset, comprising 30,000 recordings of spoken digits in English collected from 60 speakers [60]. The training data collection process involved playing each audio sample while simultaneously recording mouse sensor data (e.g., timestamps and movements). These sensor data were then aligned with their corresponding audio samples and ground-truth keywords.
@@ VI-A Evaluation Tasks and Metrics
  • Short-Time Objective Intelligibility (STOI): For measuring the intelligibility of reconstructed speech using a Discrete Cosine Transform [61].
  • Perceptual Evaluation of Speech Quality (PESQ): For correlating the quality of reconstructed speech with estimated human perception [62],
  • Scale-Invariant Signal-to-Noise Ratio (SNR): For evaluating the quality of the reconstructed waveform [63].
@@ VII-D2 Influence of environment noise
Related Work Sensing Modality Volume Remote Neural Filtering Sampling Rate Resolution Channel Bitrate SNR (dB) Accuracy (%) STOI
Lamphone [37] Optical 75-95dB 4kHz 24 bits 48 kbps +24dB - 0.53
Gyrophone [39] Mechanical 75dB  200Hz 16 bits 3.2kbps -  26% -
LidarPhone [38] Laser 55-75dB 1.8kHz 16 bits 28.8 kbps - 91% -
Visual Microphone [64] Video 80-110dB 2.2kHz >10M bits >22Mbps +30dB - -
mmEavesdropper [65] Radar 60-90dB - - +17dB 93% -
Ours Optical 50-80dB 8kHz 1.83 bits 14.6 kbps +19dB 61% 0.55
@@ VIII Related Works
This section outlines the similarities and differences between our approach and previous studies on sensor side-channel attacks. Table VII compares our proposed Mic-E-Mouse and related works. We observe that the experimental metrics across these studies are closely tied to both the channel’s bitrate and the intrinsic fidelity of its signal. For instance, the video-based side channel—operating at over 22 Mbps—delivers the highest SNR, whereas the mechanical channel at just 3.2 kbps records the lowest accuracy. Given its intermediate bitrate, Mic-E-Mouse naturally achieves mid-range performance in terms of the reported metrics
@@ VIII-A Sensor-based Side-Channels
Sensor-based Side-Channels are an ever-present threat in the modern digital landscape. Attacks on computer systems via compromised sensor and peripheral access have been shown to be effective at stealing privileged information [66, 14, 67, 68, 69, 70, 71, 17]. With the further integration of these technologies into consumer-facing products, the attack surface is poised to expand significantly. We note that motion sensors are a source of vulnerabilities as well, allowing for the reconstruction of stylus input through only the careful observation of sensor readings [14]. By exploiting access to sensor data in a similar way as the Mic-E-Mouse attack, the extraction of sensitive user data is feasible in much the same manner as these prior cyber-physical attack vectors.
@@
Portable speakers and headphones are another weak point in the security model, as audio has been extracted using a magnetic side-channel attack [67]. This is done by targetting in-ear headphones and headsets, where the magnetic signals from the speaker diaphragm allow for information leakage. Notably, their method requires physical access to the victim and their vulnerable audio device. In the Mic-E-Mouse attack, we describe an attack that does not require physical proximity to the victim.
Cryptographic security is not immune to these types of attacks either [66, 72, 73, 74, 75]. It has been shown that microphone signals can be used to decipher RSA keys without physical access to a victim’s computer [66]. This attack exploits the acoustic characteristics of computers in order to extract privileged data, much like the Mic-E-Mouse attack. However, the authors use a ground-truth audio signal to extract cryptographic keys, while we use collected mouse data to reconstruct a facsimile of a ground truth audio signal.
@@ IX-A Phoneme Intelligibility
Figure 13 illustrates the “speech banana” [48], covering typical speech frequencies (200Hz–2000Hz). Our overlay of the pipeline’s filtered audio waveform confirms that this range corresponds to phoneme frequency and is thus captured reliably under the Nyquist–Shannon sampling theorem. Appendix -A provides additional details.
@@ IX-C Countermeasures
Blacklisting Vulnerable Devices. Devices featuring high-performance optical sensors (e.g., Paw3395, Paw3399) can be banned by IT policies. This can be automated at the OS level (e.g., via udev rules [76]) to de-authorize high-risk USB HID devices.

### v2

@@ V-A Targeted Mouse Sensors
We focus on two state-of-the-art mice sensors currently available in consumer-grade mice – Paw3395 and Paw3399 – both produced by the manufacturer PixArt Imaging Inc. [paw3395ds, paw3399ds]. These sensors are in the Razer Viper 8KHz [razerviper] and Darmoshark M3-4KHz [darmosharkm3], respectively. An overview of the mice sensors technology used in our testing methodology is provided in Table I. We also list similar mice that are vulnerable to the same class of exploit as the Razer and Darmoshark mice in Appendix -B. These additional mice share at least one of the following attributes: (i) a Paw3395 or Paw3399 sensor or (ii) a poll rate at or above 4KHz.
@@ V-C Mouse Data Preprocessing
To address the nonuniformity in the collected mouse signals, we apply a signal processing pipeline that includes sequential resampling and filtering stages. We use a sinc-based resampling technique that allows efficient and accurate interpolation and reconstruction of fragmented, nonuniform, and noisy signals [eldarnonuniform]. This resampling stage is equivalent to applying a modified Whittaker–Shannon interpolation formula:
@@ V-D Waveform Signal Filtering
It is important to note that environmental and sensor noise are a point of concern in our data collection scheme. Therefore, we implement a custom-tuned Wiener Filter adapted to the sensor noise introduced by having a high DPI setting. We collect 10-minute segments of the mouse jitter which are used to estimate the noise profile, and we couple the noise spectrum with the true spectrum of speech which we acquire from our selected datasets (see Sec V-E3). In real life scenarios, the attacker can obtain noise recordings by exploiting one of the methods discussed in Section III-A to collect data during times deep into the night when the user is expected to not be using the computer. If such information cannot be obtained, another alternative would be to tune the Wiener filter using the power spectral density of sections with the lowest overall signal energy as these are most likely the ones where the user is silent. By applying a Wiener filter, the attacker can efficiently increase the Signal-to-Noise Ratio (SNR) of the resulting waveform [wienerfilter, chenwiener].
@@ V-E2 Speech Keyword Classification Models
This task involves the automated extraction of significant keywords from reconstructed audio waveforms. Formally, it is defined as a mapping function that translates the predicted audio waveform signal, denoted as $\hat{Y}$, into a set of key terms $K_{1},K_{2},\ldots,K_{n}$} that encapsulate critical information within the speech. We use wav2vec2[wav2vec2], designed to extract and classify keywords. Wav2vec2 is based on a transformer architecture, enabling a large temporal receptive field that spans 25 ms of audio. The Wav2vec2 model is trained on a supervised dataset, for 20 epochs each, with a batch size of 64 and Adam optimizer.
@@ V-E3 Training Datasets
The ML models incorporated in the Mic-E-Mouse pipeline are trained using English voice data from existing speech recognition datasets, including (i) the CSTR VCTK Corpus (version 0.92) dataset, which contains approximately 54 hours of human speech recorded from 110 English speakers [vctk], and (ii) the AudioMNIST dataset, comprising 30,000 recordings of spoken digits in English collected from 60 speakers [audiomnist]. The training data collection process involved playing each audio sample while simultaneously recording mouse sensor data (e.g., timestamps and movements). These sensor data were then aligned with their corresponding audio samples and ground-truth keywords.
@@ VI-A Evaluation Tasks and Metrics
  • Short-Time Objective Intelligibility (STOI): For measuring the intelligibility of reconstructed speech using a Discrete Cosine Transform [stoi].
  • Perceptual Evaluation of Speech Quality (PESQ): For correlating the quality of reconstructed speech with estimated human perception [pesq],
  • Scale-Invariant Signal-to-Noise Ratio (SNR): For evaluating the quality of the reconstructed waveform [snr].
@@ VII-D2 Influence of environment noise
Related Work Sensing Modality Volume Remote Classification Model Neural Filtering Sampling Rate Resolution Channel Bitrate SNR (dB) Accuracy (%) STOI
Lamphone [nassilamp] Optical 75-95dB - - 4kHz 24 bits 48 kbps +24dB - 0.53
Gyrophone [gyrophone] Mechanical 75dB Dynamic Time Warping -  200Hz 16 bits 3.2kbps -  26% -
LidarPhone [LidarPhone] Laser 55-75dB CNN - 1.8kHz 16 bits 28.8 kbps - 91% -
Visual Microphone [visualmicrophone] Video 80-110dB - - 2.2kHz >10M bits >22Mbps +30dB - -
mmEavesdropper [mmeavesdropper] Radar 60-90dB LeNet-5 U-Net - - - +17dB 93% -
Ours Optical 50-80dB Wave2Vec Transformer 8kHz 1.83 bits 14.6 kbps +19dB 61% 0.55
@@ VIII Related Works
This section outlines the similarities and differences between our approach and previous studies on sensor side-channel attacks. Table VII compares our proposed Mic-E-Mouse and related works. We observe that the experimental metrics across these studies are closely tied to both the channel’s bitrate and the intrinsic fidelity of its signal. For instance, the video-based side channel [visualmicrophone]—operating at over 22 Mbps—delivers the highest SNR, whereas the optical channel of Lamphone [nassilamp] with 48 kbps delivers worse SNR. The Gyrophone [gyrophone] with its mechanical channel records the lowest accuracy owing to the low bitrate at just 3.2 kbps, and the use of Dynamic Time Warping (DTW) which performs worse than ML based approaches. Given its intermediate bitrate, Mic-E-Mouse naturally achieves mid-range performance in terms of the reported metrics
@@ VIII-A Sensor-based Side-Channels
Sensor-based Side-Channels are an ever-present threat in the modern digital landscape. Attacks on computer systems via compromised sensor and peripheral access have been shown to be effective at stealing privileged information [genkinaudio, farrukhpencil, liaomagear, yangyro, baaccel, chenmilli, humilliear, kwonghdd]. With the further integration of these technologies into consumer-facing products, the attack surface is poised to expand significantly. We note that motion sensors are a source of vulnerabilities as well, allowing for the reconstruction of stylus input through only the careful observation of sensor readings [farrukhpencil]. By exploiting access to sensor data in a similar way as the Mic-E-Mouse attack, the extraction of sensitive user data is feasible in much the same manner as these prior cyber-physical attack vectors.
@@
Portable speakers and headphones are another weak point in the security model, as audio has been extracted using a magnetic side-channel attack [liaomagear]. This is done by targetting in-ear headphones and headsets, where the magnetic signals from the speaker diaphragm allow for information leakage. Notably, their method requires physical access to the victim and their vulnerable audio device. In the Mic-E-Mouse attack, we describe an attack that does not require physical proximity to the victim.
Cryptographic security is not immune to these types of attacks either [genkinaudio, liufreq, aydinseal, zhangvm, scheperswpa]. It has been shown that microphone signals can be used to decipher RSA keys without physical access to a victim’s computer [genkinaudio]. This attack exploits the acoustic characteristics of computers in order to extract privileged data, much like the Mic-E-Mouse attack. However, the authors use a ground-truth audio signal to extract cryptographic keys, while we use collected mouse data to reconstruct a facsimile of a ground truth audio signal.
In [JerryAttack], the authors presented an approach fundamentally different from our Mic-E-Mouse 444[JerryAttack] was not published at the time of this paper’s submission, our project GitHub Repo started in May 2023, and Data was public in January 2024. Their method relies on tampering with the mouse firmware to stream the raw (26x26) pixel image at 7-bit precision. These images, sampled at 3.7 kHz to reconstruct speech, require a bitrate of 17.5 Mbps, making the attack highly discoverable and of limited practical use. Furthermore, their threat model relies on an auxiliary microphone in the victim’s vicinity to perform full waveform reconstruction. Thus, Mic-E-Mouse is the first to demonstrate a practical and truly covert method, one that uniquely uses only standard user-space x-y coordinate data, which requires no physical or firmware modifications, and needs no external microphone.
@@ IX-A Phoneme Intelligibility
Figure 13 illustrates the “speech banana” [speechbanana], covering typical speech frequencies (200Hz–2000Hz). Our overlay of the pipeline’s filtered audio waveform confirms that this range corresponds to phoneme frequency and is thus captured reliably under the Nyquist–Shannon sampling theorem. Appendix -A provides additional details.
@@ IX-C Countermeasures
Blacklisting Vulnerable Devices. Devices featuring high-performance optical sensors (e.g., Paw3395, Paw3399) can be banned by IT policies. This can be automated at the OS level (e.g., via udev rules [udevman]) to de-authorize high-risk USB HID devices.
@@ XI Conclusion
-A The "Speech Banana"
We note that certain high-frequency phonemes–or perceptually distinct units of sound–such as f, s, and th are significantly less distinguishable to our human evaluators. An explanation is derived from two important concepts – (1) the Nyquist-Shannon sampling theorem [shannonfs], and (2) a mapping of common phonemes to loudness and frequency characteristics [speechbanana]. In Figure 13, we illustrate how the polling rate affects the intelligibility of selected high-frequency phonemes in the recovered audio waveform.
When a continuous-time signal is sampled at a rate less than double its highest frequency component, the resultant discrete waveform is not able to accurately represent the original continuous-time signal [shannonfs]. For a sensor with a sample rate $F_{s}$ of 8KHz, we can obtain a useful discrete-time signal with maximum frequency up to $F_{s}/2$, or 4KHz. Similarly, for a 4KHz sensor, the maximum recoverable frequency then drops to 2KHz. Notably, the f, s, and th phonemes have a characteristic frequency higher than even the Nyquist frequency for the 8KHz sensors. Moreover, these phonemes also have a below average loudness characteristic, further contributing to the low intelligibility observed by our human testers.
Figure 14: The distribution of phonemes and their associated minimum polling rate.
We also note that the overwhelming majority (91.18%) of phonemes in the English language are within the frequency range of the 8KHz sensor, and a significant portion (80.09%) is still within the frequency range of the 4KHz sensor [cmudict, phonemeanalysis, bnclist, advbiofreq, speechbanana]. These percentages are based solely from the theoretical lower bound on required sampling frequency obtained from the Nyquist-Shannon sampling theorem. In fact, we can see empirical evidence of this in Figure 13, where the filtered signal strength drops off noticeably after the Nyquist frequency. Notably, even for a mouse with a 1KHz sensor, 42.47% of phonemes are within this upper bound. The full distribution of phoneme prevalence versus required polling rate is shown in Figure 14.
-B Additional Vulnerable Mice
TABLE VIII: An expanded selection of vulnerable mice (Red: DPI & Poll Rate, Orange: Poll Rate only, Yellow: DPI only)
increasing vulnerability
Vendor/Model Sensor DPI IPS Polling Rate (Hz) Unit Price (USD) Reconstruction
Razer Viper 8KHz Paw3399 20,000 650 8,000 \$50 91.18%
Darmoshark M3 Paw3395 26,000 650 4,000 \$45 80.09%
Darmoshark M34K Paw3395 26,000 650 4,000 \$65 80.09%
DeLUX M800 Ultra Paw3395 26,000 650 4,000 \$60 80.09%
Pulsar Gaming Gears X2H Mini Paw3395 26,000 650 4,000 \$100 80.09%
Rapoo VT3S Paw3395 26,000 650 4,000 \$55 80.09%
ThundeRobot ML901 Paw3395 26,000 650 4,000 \$60 80.09%
ThundeRobot ML903 Paw3395 26,000 650 4,000 \$80 80.09%
VGN Dragonfly F1 Paw3395 26,000 650 4,000 \$35 80.09%
Zaopin Z1 Pro Paw3395 26,000 650 4,000 \$50 80.09%
G-Wolves Hati S Plus Paw3399 20,000 650 4,000 \$160 80.09%
AtomPalm Hydrogen Paw3360 12,000 250 8,000 \$100 91.18%
AtomPalm Hydrogen 2 Paw3360 12,000 250 8,000 \$100 91.18%
Zaunkoenig M2K Paw3360 12,000 250 8,000 \$350 91.18%
Darmoshark N3 Paw3395 26,000 650 1,000 \$40 42.47%
DeLUX M800 Pro Paw3395 26,000 650 1,000 \$50 42.47%
Glorious Model O 2 Paw3395 26,000 650 1,000 \$65 42.47%
Keychron M3 Paw3395 26,000 650 1,000 \$55 42.47%
Redragon M994 Paw3395 26,000 650 1,000 \$35 42.47%
Edifier Hecate G3M Pro Paw3399 20,000 650 1,000 \$70 42.47%
Razer Basilisk V2 Paw3399 20,000 650 1,000 \$40 42.47%
Razer Basilisk V3 Paw3399 20,000 650 1,000 \$80 42.47%
Razer Basilisk Ultimate Paw3399 20,000 650 1,000 \$105 42.47%
Razer DeathAdder V2 Paw3399 20,000 650 1,000 \$25 42.47%
Razer DeathAdder V2 Pro Paw3399 20,000 650 1,000 \$70 42.47%
Razer Viper Ultimate Paw3399 20,000 650 1,000 \$160 42.47%
Figure 15: Comparison of the sweep response for the three selected membranes.
A list of similar mice that are also vulnerable to the same class of exploit as the Razer and Darmoshark is presented in Table VIII. These additional mice share at least one of the following attributes: (1) a Paw3395 or Paw3399 sensor or (2) a polling rate at or above 4KHz. We list sensors with a 4KHz or 8KHz. Table VIII is ordered first by sensor vulnerability exposure and then alphabetically. Note that the asterisk in the last column of the table represents the theoretical upper bound on reconstruction percentage as shown in Appendix -A. This reconstruction percentage is based on the polling rate of the sensor, whereas the overall vulnerability rating takes into account multiple other factors such as DPI, IPS, etc.
One important consideration is that certain mouse vendors market the sensors by using brand-specific names. An example of this is Razer using the brand name Focus+ in lieu of the actual sensor used– he Paw3399. A similar case applies to Glorious Gaming’s BAMF2.0, which is a derivative of the Paw3395 sensor. We consider this kind of sensors as derivatives of the original PixArt Imaging sensors in our analysis and discussion. Table VIII shows a list of 26 unique products using either the Paw3395 or Paw3399 sensors sorted according to their potential susceptibility to our attack.
We note the large size of the computer hardware market validates the growing ubiquity of these high-performance mice with state-of-the-art optical sensors. For example, Razer had total hardware sales of US\$1,452.4 million during Fiscal Year 2021 [razer21report], including desktops, laptops, and other computing equipment. For further evidence of the growing market demand, it’s anticipated that the global market for computer mice will increase to US\$2.61 billion by the end of 2023, with a Compounded Annual Growth Rate (CAGR) of 8.1% [mice23demand]. Then, it’s clear that this type of high-performance mouse sensors will become more pervasive over time.
-C Full Membrane Study
We present a sweep response for various types of membranes, including paper, plastic, and cardboard. These are a representative sample of common types of surfaces a mouse may be placed on in a real-world environment. The sweep response, along with a comparison to the ground truth signal is presented in Figure 15. The different responses are mainly attributed to three properties:
  • The thickness $W$ of the material significantly influences acoustic absorption and reflection. Thicker materials exhibit high sound absorption, leading to a dampened sweep response, whereas thinner materials may reflect sound more efficiently, resulting in an accentuated response.
  • The density $\rho$ of the material dictates both the energy needed to vibrate the surface, as well as the momentum it carries before the signal dies down. Lighter materials, such as paper, are mismatched with the ambient noise and only acquire the injected sounds.
  • The speed of sound in the material $\nu$. Materials with a higher $\nu$ facilitate rapid sound transmission, potentially yielding a less distorted response. In contrast, materials with a lower $\nu$ can delay wave propagation, altering the phase and amplitude response in the sweep.
Modeling the complete response given material properties is not trivial, given the existence of standing waves, non-linear responses, elastic attenuation, and potentially other physical phenomenon. However, we notice that these differences are mainly translated in two aspects in the extracted signals:
  • Harmonics: The emergence of 2nd and 3rd order harmonics in plastic and paper membranes suggests signal degradation. Nonetheless, this redundancy in frequency acquisition offers a compensatory mechanism, providing machine learning models with a ’second chance’ to discern the underlying signal.
  • The noise profile: The susceptibility to random noise varies significantly across materials. Notably, the paper membrane displays a distinct response, characterized by its ’quietness’ in the sweep. This is attributed to a minimal presence of standing waves and rapid energy dissipation, owing to paper’s low momentum energy, in contrast to cardboard and plastic.